OpenStudy (safia21):

Write the equation of the line parallel to x = -16 and passes through the point (9, -10).

OpenStudy (anonymous):

parallel=same slope

OpenStudy (anonymous):

y=-16x+b

OpenStudy (anonymous):

-10=-16(9)+b. Solve for b

OpenStudy (anonymous):

-10=-144+b 134=b

jimthompson5910 (jim_thompson5910):

equation is x = -16 gandalfwiz not y = -16

OpenStudy (anonymous):

:p me stupid

jimthompson5910 (jim_thompson5910):

that's ok, the answer is actually x = 9 since this is also a vertical line but it passes through (9,-10)

OpenStudy (safia21):

how do you do this one? Write the equation of the line perpendicular to 9x + 2y = 18 and passes through the point (1, 4).

OpenStudy (anonymous):

That's when you do it my way :). get your first equation into y=mx+b form, then follow my steps above

OpenStudy (anonymous):

9x+2y=18 2y=-9x+18 y=-4.5x+9

OpenStudy (anonymous):

perpendicular= negative reciprocal slope

OpenStudy (anonymous):

The negative reciprocal is the number flipped upside down and multiplied by a negative.

OpenStudy (anonymous):

-4 1/2=-9/2 2/9

OpenStudy (anonymous):

Now u have y=2/9x+b

OpenStudy (anonymous):

plug in your point for the y and x

OpenStudy (anonymous):

4=2/9(1)+b 4=2/9+b 4-2/9=b

OpenStudy (anonymous):

Since 4- 2/9 gives you tons of decimals, I'd leave it like that. Your equation is y=2/9x+(4- 2/9)

OpenStudy (safia21):

a. 9x - 2y = 1 b. 9x + 2y = 17 c. 2x - 9y = -34 d. 2x + 9y = 38

OpenStudy (safia21):

which one?

jimthompson5910 (jim_thompson5910):

of course, you can combine 4-2/9 to get 34/9

jimthompson5910 (jim_thompson5910):

choice C

jimthompson5910 (jim_thompson5910):

another way to do it is swap the x and y coefficients and change the sign. From there, plug in x=1 and y=4 to find the right hand side

OpenStudy (safia21):

can you start from the beginning please i am so confused

jimthompson5910 (jim_thompson5910):

sure, one sec

OpenStudy (safia21):

thanks :)

jimthompson5910 (jim_thompson5910):

start with 9x + 2y = 18 ANY line perpendicular to this will be of the form 2x-9y = C where C is any number All I did here was swap the x and y coefficients 9 and 2 and then I changed the sign of the final y coefficient

jimthompson5910 (jim_thompson5910):

So we now have 2x-9y = C Since this perpendicular line must go through (1,4), this means that x=1 and y=4 Plug these values in the equation above to get 2x-9y = C 2(1)-9(4) = C 2-36 = C -34 = C C = -34 So the equation of the perpendicular line that passes through (1,4) is 2x-9y = -34

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (safia21):

where did you get 1 and 4

jimthompson5910 (jim_thompson5910):

from the point (1,4)

jimthompson5910 (jim_thompson5910):

that the perpendicular line must pass through

OpenStudy (safia21):

o okay got it thanks!

jimthompson5910 (jim_thompson5910):

great, np